If a right triangle has a hypotenuse of 20 units and a leg of 16 units, what is the length of the other leg?

Answer by Zucchini(70)

Use the equation a^2 + b^2 = c^2
A is a leg, b is the other leg, c is the hypotenuse.
Plug in what you know:
16^2 + b^2 = 20^2
256 + b^2 = 400
B^2 = 144
B = 12
The length of the other leg is 12 units